# On Weighted Arboricity: Conductance–Resistance Bounds and Monoid Structure

### v3

@@ Abstract
We study a conductance-weighted arboricity for a finite simple undirected graph $G=(V,E,c)$ with a conductance assignment $c:E\to[0,\infty)$:
@@
This functional reduces to the classical Nash–Williams identity when $c\equiv 1$, is isomorphism invariant, monotone under subgraphs and edge additions, positively homogeneous, and convex. We prove sharp global bounds
@@ 1 Introduction
The arboricity of a graph $G$, defined as the minimum number of forests whose union covers every edge of $G$, admits a classic optimization characterization due to Nash–Williams [1]: the arboricity is the ceiling of the maximum, over all connected subgraphs $H$ with $|V(H)|\geq 2$, of |E(H)|/(|V(H)|1)|E(H)|/(|V(H)|-1). This density interpretation has proven to be useful in structural and algorithmic graph theory [2, 3, 4].
Weighted analogues of arboricity have been studied in several directions. When edges carry positive integer weights, Gabow and Westermann [2] gave min-max theorems for the minimum number of forests needed to cover each edge at least its weight. Real-weighted versions, especially with weights interpreted as capacities or costs, appear in multicommodity flow and congestion minimization [5]. However, the natural conductance-weighted analogue, where the density numerator is the sum of conductances (reciprocal resistances) rather than cardinality or integer weights, has received surprisingly little direct attention.
Independently, the language of electrical networks has become a standard tool for studying graph densities. Foster’s theorem [6] and Rayleigh’s monotonicity principle imply that the effective resistance $R^{(G)}_{\text{ eff }(e)}$ between any two vertices is at most the reciprocal of the number of edge-disjoint paths, and this perspective has been used to bound spanning-tree densities and Laplacian matrix eigenvalues [7, 8]. Despite the close formal analogy between conductance-weighted spanning-tree counts and effective resistance, no systematic study of the resulting functional appears in the literature.
The purpose of this article is to fill this gap. For a finite undirected graph $G$ equipped with nonnegative edge conductances $c:E\to[0,\infty)$, we define the conductance-weighted arboricity
@@
    The functional $\mathcal{A}_{c}(G)$ inherits all expected analytic properties from the classical case: isomorphism invariance, monotonicity under edge addition and weight increase, positive homogeneity, convexity, and sharp global bounds $\max c(e)\leq\mathcal{A}_{c}(G)\leq\sum c(e)$, with attainment (Section 3).
@@ 3.2 Basic properties
Proposition 3.3 (Reduction to the classical density).
@@
$\mathcal{A}_{c}(G)\;=\;\max_{H\subseteq G\text{ conn}}\frac{|E(H)|}{|V(H)|-1},$
the Nash–Williams density. In particular, $\lceil\mathcal{A}_{c}(G)\rceil$ equals the (unweighted) arboricity.
@@
(ii) For any $e=\{u,v\}$, the subgraph on $\{u,v\}$ with edge $e$ gives $D_{c}(H)=c(e)$; taking $e$ with maximal weight gives the bound.
(iii) For any spanning tree $T$ of $G$, $\mathcal{A}_{c}(G)\geq\big(\sum_{e\in E(T)}c(e)\big)/(|V(G)|-1)\geq\underline{c}$.
(iv) There are finitely many connected subgraphs of $G$; thus $\{D_{c}(H)\}$ has a maximum. ∎
@@
[…]
Figure 1: Two weighted trees with edge conductances labeled. Since every connected subgraph of a tree is a subtree, the density $D_{c}(H)$ is an average of edge conductances over $H$, hence maximized by a single edge of largest conductance; therefore $\mathcal{A}_{c}(T)=\max_{e\in E(T)}c(e)$.
@@ 4.4 Separation from totals or degrees
This is, of course, a somewhat trivial implication of the reduction to classical arboricity for $\mathcal{A}_{1}(G)$. Further results on heterogeneous $c$ are discussed in Section 6.
@@ 7 Conclusion
together with the associated extremal quantity obtained by maximizing over connected subgraphs. The guiding theme is that embedding conductances into the Nash–Williams framework produces an invariant that is both structurally similar to classical arboricity and also naturally tethered to effective resistance and electrical network geometry. On the theoretical side, we developed a usable upper bound for the extremal density in terms of effective resistances in the ambient graph, which is explicit once effective resistances are available. The resulting inequality behaves well under scaling, interfaces cleanly with classical families, and is computable in practice via standard linear algebra. A brief computational exhibit complements the theory. Using Python, we generated overlay and tightness-ratio plots comparing the baseline density to the explicit certificate across graph sizes. We emphasized the hypercube family $Q_{d}$, where edge–transitivity implies that all edges share the same effective resistance, so $\mathrm{CSBound}(H{=}G)$ can be evaluated from a single grounded–Laplacian solve per dimension. Both in the unit-conductance base case and under random two-point conductances, the certificate tracks the observed densities closely, and the tightness ratio trends toward $1$ as $d$ grows, supporting the view that the bound is sharp on highly symmetric families.

### v4

# Extensions of Real-Weighted Fractional Arboricity: Conductance–Resistance Bounds and Monoid Structure

@@ Abstract
We study a conductance-weighted arboricity for a finite simple undirected graph $G=(V,E,c)$ with a conductance assignment $c:E\to(0,\infty)$:
@@
This functional reduces to the fractional arboricity $\gamma(G)$ when $c\equiv 1$, is isomorphism invariant, monotone under subgraphs and edge additions, positively homogeneous, and convex. We prove sharp global bounds
@@ 1 Introduction
The arboricity of a graph $G$, defined as the minimum number of forests whose union covers every edge of $G$, admits a classic optimization characterization due to Nash–Williams [1]: the arboricity is the ceiling of the maximum, over all connected subgraphs $H$ with $|V(H)|\geq 2$, of |E(H)|/(|V(H)|1)|E(H)|/(|V(H)|-1). The fractional arobricity, denoted $\gamma(G)$, is the real-valued analogue resulting from the simple omission of the ceiling function from the classical case; this density interpretation has proven to be useful in structural and algorithmic graph theory [2, 3, 4].
Weighted analogues of arboricity have been studied in several directions. When edges carry positive integer weights, Gabow and Westermann [2] gave min-max theorems for the minimum number of forests needed to cover each edge at least its weight. Real-weighted versions, especially with weights interpreted as capacities or costs, appear in multicommodity flow and congestion minimization [5]. However, the natural conductance-weighted analogue, where the density numerator is the sum of conductances (reciprocal resistances) rather than cardinality or integer weights, has received surprisingly little direct attention. Many intended applications of this paper benefit from real-valued parameters (such as covariates in regressions), so we intentionally drop the ceiling in Nash–Williams’ definition and move from integer- to real-valued weights, using conductances due to their many useful qualities; namely, the language of electrical networks has become a standard tool for studying graph densities. Foster’s theorem [6] and Rayleigh’s monotonicity principle imply that the effective resistance $R^{(G)}_{\text{ eff }(e)}$ between any two vertices is at most the reciprocal of the number of edge-disjoint paths, and this perspective has been used to bound spanning-tree densities and Laplacian matrix eigenvalues [7, 8].
Despite the close formal analogy between conductance-weighted spanning-tree quantities and effective resistance, few systematic studies relating arboricity to effective resistance exist within the literature. Most similar to this paper, effective resistances have been used to bound fractional arboricity in the unweighted setting, and in fact a bound which takes similar structure to our weighted arboricity functional is procured from use of Rayleigh monotonicity in [9]. This research diverges from their bound on fractional arboricity in that it studies a fundamentally different object, a real-weighted functional with exogenous edge weight assignments, and concerns itself on proving various properties about that functional rather than using electrical network theory to bound the unweighted fractional arboricity functional; in particular, we derive a local Foster inequality for arbitrary conductance assignments and a resistance bound, and we complement the theory with computational experiments under heterogeneous conductance sampling.
The purpose of this article is to continue to fill this gap between arboricity and electrical network theory due to their surprisingly numerous similarities. For a finite undirected graph $G$ equipped with nonnegative edge conductances $c:E\to(0,\infty)$, we define the conductance-weighted arboricity
@@
    The functional $\mathcal{A}_{c}(G)$ inherits all expected analytic properties from the fractional case: isomorphism invariance, monotonicity under edge addition and weight increase, positive homogeneity, convexity, and sharp global bounds $\max c(e)\leq\mathcal{A}_{c}(G)\leq\sum c(e)$, with attainment (Section 3).
@@ 3.1 Setup
Remark 3.3 (Exclusion of zero edge weights).
Readers may notice that 0 was intentionally excluded from the definition of conductance assignment; this is because for the purposes of this research, such an edge weight will contribute nothing to the sum in the numerator of $D_{c}(G)$. Indeed, excluding $c=0$ edge weights may be interpreted as calculating maxima over connected components of the induced subgraph $G^{\prime}=(V,\{e\in E(G):c(e)>0\})$, where trivially $\mathcal{A}_{c}(G)=\mathcal{A}_{c}(G^{\prime})$, and thus we are left without a reason to include them in the definition of.
@@ 3.2 Basic properties
Proposition 3.4 (Reduction to fractional arboricity).
@@
$\mathcal{A}_{c}(G)\;=\;\max_{H\subseteq G\text{ conn}}\frac{|E(H)|}{|V(H)|-1}=\gamma(G),$
the fractional arboricity. In particular, $\lceil\mathcal{A}_{c}(G)\rceil$ equals the classical (unweighted) Nash–Williams arboricity.
@@
(ii) For any spanning tree $T$ of $G$, $\mathcal{A}_{c}(G)\geq\big(\sum_{e\in E(T)}c(e)\big)/(|V(G)|-1)\geq\underline{c}$.
(iii) There are finitely many connected subgraphs of $G$; thus $\{D_{c}(H)\}$ has a maximum. ∎
Corollary 3.10 (Weighted arboricity of trees).
Let T be a tree with a conductance assignment. Then $\mathcal{A}_{c}(G)=\max_{e\in E(T)}c(e).$
Proof.
Since every connected subgraph of a tree is a subtree, the density $D_{c}(H)$ is an average of edge conductances over $H$, hence maximized by a single edge of largest conductance; therefore $\mathcal{A}_{c}(T)=\max_{e\in E(T)}c(e)$. ∎
[…]
Figure 1: Two weighted trees with edge conductances labeled.
@@ 4.4 Separation from totals or degrees
This is, of course, a somewhat trivial implication of the reduction to fractional arboricity for $\mathcal{A}_{1}(G)$. Further results on heterogeneous $c$ are discussed in Section 6.
@@ 7 Conclusion
together with the associated extremal quantity obtained by maximizing over connected subgraphs. The guiding theme is that embedding conductances into the unweighted arboricity framework produces an invariant that is both structurally similar to fractional arboricity and also naturally tethered to effective resistance and electrical network geometry. On the theoretical side, we developed a usable upper bound for the extremal density in terms of effective resistances in the ambient graph, which is explicit once effective resistances are available. The resulting inequality behaves well under scaling, operates cleanly with classical families, and is computable in practice via standard linear algebra. A brief computational exhibit complements the theory. Using Python, we generated overlay and tightness-ratio plots comparing the baseline density to the explicit certificate across graph sizes. We emphasized the hypercube family $Q_{d}$, where edge–transitivity implies that all edges share the same effective resistance, so $\mathrm{CSBound}(H{=}G)$ can be evaluated from a single grounded–Laplacian solve per dimension. Both in the all-ones base case and under random two-point conductances, the certificate tracks the observed densities closely, and the tightness ratio trends toward $1$ as $d$ grows, supporting the view that the bound is sharp on highly symmetric families.